Find the polynomial equation over R of lowest degree which is satisfied by 1− i
and 3 + .i2
1
Expert's answer
2019-03-28T04:14:44-0400

P(x) = (x-1+i)(x-1-i)(x-3-2i)(x-3+2i) = (x2 - 2x + 2)(x2 - 6x + 13) = x4 - 8x3 + 27x2 - 38x +26;


Answer: P(x) = x4 - 8x3 + 27x2 - 38x +26;